Please answer the above with these given numbers
Given the following MW
P-aminophenol - 109.13 g/mol
Acetic anhydride - 102.09 g/mol
Acetaminophen - 151.16 g/mol
Acetic acid - 60.05 g/mol

Thank youuuu

media%2Fcf4%2Fcf465b41-1bb7-4d64-ac10-dc
Answer #1

a)

Number of moles of p-aminophenol = mass of p-aminophenol used / molecular mass

                                                          = 0.150 / 109.13

                                                           = 1.37 mmol

b)

Density of acetic anhydride = 1.08 g/cm3

Mass of acetic anhydride used = volume * density

                                                 = 0.3 mL * 1.08 g/cm3

                                                 = 0.324 g

Number of moles of acetic anhydride used = mass of acetic anhydride used / molecular mass

                                                                    = 0.324 / 102.09

                                                                    = 3.17 mmol

c) Based on the above calculations it is clear that p-aminophenol is the limiting reagent. Limiting reagent in a chemical reaction is the reagent which consumes completely and thus limits formation of product. Here acetic anhydride is used as excess amount and it does not uses completely in this reaction.

d)

The theoretical yield means 100% yield. 100% yield means complete conversion of starting material into product.

Thus the theoretical yield of this reaction is 1.37 mmol product.

Mass of the product = number of mmol * molecular mass

                                = 1.37 * 151.16

                                = 0.207 g

e)

Yield = (mass of product you got / mass of the product if 100% yield ) * 100

          = (0.180 / 0.207 ) * 100

          = 87 %